OpenStudy (anonymous):

suppose you invest $1600 at an annual rate of 4.6% compounded continuously. How much will you have in the account after 4 years?

OpenStudy (anonymous):

$1,894.40

OpenStudy (anonymous):

I=PRT P = 1,600 R =0.046 T = 4 1,600 * 0.046 * 4 = 73.6 * 4 = 294.4 The answer is 294.4 ;)